OpenStudy (anonymous):

One side of an isosceles triangle whose perimeter is 42 units measures 10 units. FInd the area of the triangle.

OpenStudy (anonymous):

from what i've read: isoceles triangles are triangles with two congruent sides and two congruent angles. but i can't figure what side is that 10 units does it belong with the two congruent sides or the other side...

OpenStudy (wolf1728):

Perimeter of isosceles triangle (2*leg) + base The measure of 10 could be a leg or a base If 10 is a leg then both legs = 20 and the base would be 22 This cannot be a triangle because the longest side of a triangle must be less than the sum of the other 2 sides. (22 is > 20 which eliminates this from occurring.) So the 10 length must be the base and each leg is 16.

OpenStudy (anonymous):

then i can use A=1/2bh ? thus obtaining the answer?

OpenStudy (wolf1728):

I'm working on the area right now

OpenStudy (anonymous):

i got the altitude =\[\sqrt{231}\] and the area = \[5\sqrt{231}\] please check thanks for helping...

OpenStudy (wolf1728):

The area of a triangle = (base * height) / 2 base / 2 = 5 height² = 16² - 5² height = sq root(231) height = 15.1986841536 area = (15.1986841536 *5) area = 75.9934207679

OpenStudy (wolf1728):

Looks as if we agree :-)

OpenStudy (anonymous):

\[5\sqrt{231}= 75.99342077\] close answer i used the formula for getting the altitude which involves that semi perimeter.. and used it for getting the area. thanks for helping.. (gives medal) :))

OpenStudy (wolf1728):

thanks